OpenStudy (anonymous):

solve for a for the equation: sin(2a)+3=4.

OpenStudy (anonymous):

sin(2a) = 4-3 What gives a value of sine of one?

myininaya (myininaya):

sin(2a)=1 remember sin(pi/2)=1 sin(pi/2+2npi)=1 for n=...-2,-1,0,1,2,3,... so we need to solve 2a=pi/2+2npi for a

myininaya (myininaya):

mutiply both sides by 1/2 an you are done!